OpenStudy (anonymous):

3.789*46.34*3.2/0.2*2*0.3

OpenStudy (anonymous):

516.863232/.12=4307.1936

OpenStudy (jmark):

3.789*46.34*3.2/0.2*2*0.3 = (561.863232)/ (0.12) = 4682.1936 ICSE question papers for class 9 - http://icse.edurite.com/icse-question-papers/icse-question-papers-for-class-9.html